p-3q = sinx+cosy


1
Expert's answer
2021-07-14T18:17:27-0400
"p-3q = \sin x+\cos y"

"p- \sin x=3q +\cos y=a (say),"

where "a" is an arbitrary constant.

We have


"\begin{cases}
 p-\sin x=a \\
 3q+\cos y=a
\end{cases}=> =\begin{cases}
 \dfrac{\partial z}{\partial x}=a+\sin x \\
 \dfrac{\partial z}{\partial y}=\dfrac{1}{3}(a-\cos y)
\end{cases}"

"dz=(a+\sin x)dx+(\dfrac{1}{3}(a-\cos y))dy"

"z=ax-\cos x+\dfrac{1}{3}ay-\dfrac{1}{3}\sin y-b"

Or


"z=a(x-\dfrac{1}{3}y)-(\cos x+\dfrac{1}{3}\sin y+b)"